OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4. y = 3x - 12 y = -3x y = -3x + 12 y = 3x

OpenStudy (anonymous):

hey parallel means what?

hero (hero):

@ivandelgado, where exactly are you stuck in this problem?

OpenStudy (anonymous):

i dont no the answer?

hero (hero):

You mean, you don't know the first step to solve it.

OpenStudy (anonymous):

@ivandelgado parallel means they have exactly the same slope. look at the chocices

hero (hero):

@madgrieve, you assume he knows what slope means.

OpenStudy (anonymous):

y = 3x ?

hero (hero):

@ivandelgado, guessing isn't going to get you anywhere

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Since the line is parallel to y=-3x-4, our line will have the same slope as this one, -3x. Remember, all parallel lines have the same slope. For this problem, you have to know the point-slope formula. y-y1=m(x-x1). The m is the slope in this formula. the x1 and the y1 is the coord. For this problem, our (x1,y1) is (2,6). Also, our m for this problem will be -3, since all parallel lines have the same slope. Now since we know what everything is and how to do it we'll solve it. y-6=-3(x-2) First we distribute this. y-6=-3x+6 Now we add 6 on both sides. y=-3x+12 is what we get, and this is our answer :)

OpenStudy (precal):

 m is your slope value parallel means they are the same